# Vision Beyond Boundaries: An Initial Design Space of Domain-specific Large Vision Models in Human-robot Interaction

### v1

@@ 4.2. Data Analysis
During our brief interview sessions, we collected some valuable insights. The majority of participants expressed positivity of the design space and its three dimensions. One participant, whose research interests align closely with the proposed topic, mentioned, ”The design space you’ve introduced seems like it will be a useful tool and guide for my future research endeavors to a significant extent, as it clearly lists the essential components.” In terms of the three dimensions, another participant appreciated the strategy of outlining specific domains of focus before diving into practical research challenges brought by LVMs: ”This approach of summarizing key domains before tackling practical research problems can streamline the process of identifying the precise area of interest for applying the vision model. It not only saves time but also leads to enhanced outcomes.” Moreover, a participant with years of experience in computer vision and robotics commented on the utility of the HRI contextual categorization, stating, ”This categorization of contexts will aid in laying down the essential structure for the interactive system I plan to develop in the future, especially since my prior experience in this area is limited.”
@@ 5. Discussion
We believe that our design space offers a pioneering framework for designing future HRI architectures, utilizing LVMs across specific domains. In this section, we delve into various facets of the proposed design space.
@@
The ideation is based on the existing work and emerging trends. While numerous studies have focused on integrating various vision models in HRI, primarily for enhanced visual detection and robot control, there is a lack of structured approaches in utilizing LVMs for efficient interaction between humans and robots. Given the current surge in LLMs, we anticipate a similar revolutionary trajectory for LVMs in following years. The concept of domain-specific LVMs, first promoted and advocated by Landing AI 555https://landing.ai/, has shown superior performance over traditional models. This underscores the need for a foundational guideline that assists in designing future HRI systems, delineating specific contexts, vision tasks, and domains.

### v2

@@ 4.3. Summary of Qualitative Feedback
The majority of participants expressed positivity of the design space and its three dimensions. One participant, whose research interests align closely with the proposed topic, mentioned, ”The design space you’ve introduced seems like it will be a useful tool and guide for my future research endeavors to a significant extent, as it clearly lists the essential components.” In terms of the three dimensions, another participant appreciated the strategy of outlining specific domains of focus before diving into practical research challenges brought by LVMs: ”This approach of summarizing key domains before tackling practical research problems can streamline the process of identifying the precise area of interest for applying the vision model. It not only saves time but also leads to enhanced outcomes.” Moreover, a participant with years of experience in computer vision and robotics commented on the utility of the HRI contextual categorization, stating, ”This categorization of contexts will aid in laying down the essential structure for the interactive system I plan to develop in the future, especially since my prior experience in this area is limited.”
@@ 5. Discussion
In this section, we delve into various facets of the proposed design space.
@@
The ideation is based on the existing work and emerging trends. While numerous studies have focused on integrating various vision models in HRI, primarily for enhanced visual detection and robot control, there is a lack of structured approaches in utilizing LVMs for efficient interaction between humans and robots. Given the current surge in LLMs, we anticipate a similar revolutionary trajectory for LVMs in following years. The concept of domain-specific LVMs, first promoted and advocated by Landing AI, has shown superior performance over traditional models. This underscores the need for a foundational guideline that assists in designing future HRI systems, delineating specific contexts, vision tasks, and domains.